OpenStudy (anonymous):

How long will it take a $500 investment to be worth $1000 if it is continuously compounded at 12% per year? (Give the answer to two decimal places.)

OpenStudy (anonymous):

 A= 1000 P=500 r=12% or 0.12 n = 1 1000 = 500(1+0.12)^t 1000/500 = 1.12^t 2=1.2^t = log1.12 (2) = 6.12 years or 6 years